# Perceptual compression

### v1

@@ 4.1 Experimental setup
Implementation details. We base our model off a text-conditioned latent diffusion model, with latent space provided by a convolutional autoencoder and diffusion model with an architecture similar to GLIDE (Nichol et al., 2022). The model is pre-trained on a proprietary dataset consisting of around 300M image-caption pairs. We use BLIP-2 (Li et al., 2023) to obtain image captions for conditioning the decoder, and limit the maximum caption length to 32 tokens. The hyper encoder consists of nine residual convolutional blocks, and includes a number of downsampling layers depending on the spatial map resolution that is selected to obtain a target bitrate. In total, the hyper encoder contains between 4M and 8M parameters. During all our experiments, the autoencoder weights are frozen. We train the hyper encoder and finetune the diffusion model on OpenImages (Kuznetsova et al., 2020), similar to Lei et al. (2023) and Muckley et al. (2023). We use random $512\!\times\!512$ crops, and instead of finetuning the full diffusion we found it beneficial to only finetune the linear layers present of the diffusion model, representing around 15% of all weights.
Datasets. For evaluation, we use the Kodak dataset (kod, ) as well as MS-COCO 30k which consists of 30,000 images. On COCO we evaluate at resolution $256\!\times\!256$ by selecting the same images from the 2014 validation set (Lin et al., 2014) as Hoogeboom et al. (2023) and Agustsson et al. (2023).
We evaluate at resolution $512\!\times\!512$ on the 2017 training set (Caesar et al., 2018), which is the same resolution used for evaluation by Lei et al. (2023), and use captions and label maps for some metrics.
@@
Figure 4: Evaluation on COCO at $256\!\times\!256$ res.
Metrics. To quantify image quality we use FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018), which match feature distributions between sets of original images and their reconstructions. We measure distortion using MS-SSIM (Wang et al., 2003) and LPIPS (Zhang et al., 2018). In Sec. B.1 we also report PSNR, but find these distortion metrics to be less meaningful for low rates. We use the COCO annotations to evaluate preservation of semantics. We compute the CLIP score to measure global alignment of reconstructed samples with ground truth captions (higher values are better) (Hessel et al., 2021). Unless otherwise specified, we condition our model on captions generated by BLIP2, which are different from the ground-truth captions used to compute the CLIP score. Second, we adopt the mean Intersection over Union (mIoU) metric to measure the local semantic preservation, see e.g. Sushko et al. (2022). In this case, we pass reconstructed images through a pre-trained semantic segmentation model, and compare the predicted semantic maps to the ground truth ones. We take the pretrained ViT-Adapter segmentation network (Chen et al., 2023), which was trained on the COCO train split. For Kodak we only report LPIPS and MS-SSIM, as the dataset is too small (24 images) to reliably compute FID and KID, and does not come with semantic annotations.
Baselines. We include representative codecs from several families. VTM is a state-of-the-art hand-crafted codec used in the VVC video codec (vtm, ). Among neural compressors using adversarial losses, we include MS-ILLM (Muckley et al., 2023) which improves over HiFiC (Mentzer et al., 2020) using an adversarial loss based on multi-class classification. For HiFiC we include their model with the lowest available bitrate, and add MS-PatchGAN, which is a close reproduction of HiFiC by Muckley et al. (2023). We compare to multi-realism approach of Agustsson et al. (2023), at their lowest bitrate with the best realism. Text-Sketch (Lei et al., 2023) is the only prior work we are aware of that evaluates image compression at bitrates below 0.01, it is based on text-conditioned diffusion models. We include two other recent diffusion-based approaches: DIRAC (Ghouse et al., 2023) and HFD/DDPM (Hoogeboom et al., 2023). For details on replication of baselines see App. A.
@@ 4.2 Main results
Comparison to state-of-the-art methods. In Fig. 3 we compare our results to state-of-the-art codecs. We observe that PerCo (ours) yields significantly lower (better) FID and KID compared to other approaches at lower bitrates ($<$0.04), and that our FID and KID curves are much flatter, indicating a decoupling of realism and bitrate that is not seen for other methods.
For the semantics-related CLIP and mIoU metrics we also obtain consistent improvements over all other methods, in particular at low rates. For LPIPS, PerCo is better than Text-Sketch and VTM, but somewhat worse than other approaches for rates $>$0.01 bpp. Similarly, for MS-SSIM we improve over Text-Sketch, but are worse than other methods. However, as we illustrate in Figs. 10 and 14 in Sec. B.2, such similarity metrics are not necessarily meaningful at low bitrates, as they do not capture realism.
In Fig. 4 we compare to HFD/DDPM (using their two lowest bitrates, 0.107 and 0.175 bpp) on FID on MS-COCO 30k at $256\!\times\!256$ resolution, as used in their paper. In the same figure we also include MS-ILLM and the multi-realism approach of Agustsson et al. (2023). We find that PerCo yields best FID scores across the tested bitrates.
@@
Qualitative comparisons. We provide qualitative comparison of PerCo to Text-Sketch, VTM and MS-ILLM in Fig. 1. We observe that PerCo yields reconstructions that are of higher quality and that preserve content and appearance of the originals better. VTM and MS-ILLM reconstructions are blurry and/or blocky, while Text-Sketch strongly deviates from the originals in color.
In Fig. 5, we show two more samples with additional zooms. Again, PerCo yields reconstructions with less artefacts, more faithful colors, and realistic details (see zooms). Here, when using text-only conditioning for Text-Sketch (PIC, 0.0025 bpp), it is not able to achieve spatial alignment nor overall color preservation, due to the lack of local image encoding.
4.3 Ablations
@@
Diversity in reconstructions. With our model, reconstructions are sampled from a conditional diffusion model. Ideally, the conditional distribution concentrates all mass perfectly on the original input image. In practice, this does not happen due to limited model capacity, training data, and bitrate. In particular, we expect that at lower bitrates —where there is less information about the original sample— the conditional distribution would have more variance, and we expect to observe more diversity among the samples. The reconstruction at two bitrates in Fig. 6 confirm this tendency.
Conditioning modalities. In Fig. 7, we evaluate PerCo using each conditioning modality separately (textual or spatial), to analyze the contribution of each. We compare BLIP-2 (Li et al., 2023) to generate captions (default) to using IDEFICS (Laurançon et al., 2023) which produces more detailed descriptions, and ground-truth (GT) captions. When not using descriptive captions, we consider conditioning on a learned textual embedding that is constant across images. We also tried using an empty text or on a fixed generic text —“A high quality photograph.”— for all images, and observed similar or slightly worse performance as the learned text embedding setting.
@@
The model using the constant text embedding (PerCo w/ only Spatial Encoding) performs significantly worse than the other models in terms of FID and CLIP score. For the LPIPS and mIoU metrics, which rely on the alignment of the original and reconstructed images, the results are comparable to other models. This can be attributed to the fact that accurate alignment is primarily achieved through spatial feature conditioning, which remains unchanged in this model. Consistently, we also notice that when using textual condition only (blue, orange, and red square), the performance in terms of LPIPS and mIoU is worse than for the models that use spatial conditioning, while the CLIP scores are in line with the models using both textual and spatial conditioning. Both types of conditioning contribute to improving FID, and the text-only conditioned models perform in between the models with only spatial conditioning and those with both types of conditioning.
Between models using captions from BLIP (blue curves) and IDEFICS (orange curves) in combination with spatial conditioning, we observe similar FID scores and somewhat better CLIP scores for the more detailed IDEFICS captions. In terms of LPIPS and mIoU we find that the models with BLIP captions perform better, which is largely due to bitrate required to encode the longer captions: leaving out the textual encoding cost, both models perform very similarly. Overall, using ground-truth captions (GT, red curves) yields similar results as using BLIP, while BLIP and IDEFICS yield slightly better mIoU and CLIP scores. When conditioning on ground-truth captions, we take a different caption than the one used to compute CLIP scores, this is possible as COCO contains five captions per image.
@@
Classifier-free guidance. In Tab. 1 (left), we compare three options regarding classifier-free guidance (CFG) in our model (i) not using CFG, (ii) applying CFG on the text and spatial feature maps, i.e. contrasting full conditioning to no conditioning (we use an constant learned text embedding, and fixed all-zeros feature maps for spatial conditioning), and (iii) applying CFG to text-conditioning only, i.e. always conditioning on the spatial feature map. We find that best results are obtained when CFG is applied to textual conditioning only, which yields significantly better FID and mIoU metrics.
Impact of autoencoder and quantization. To assess to which extent performance is bounded by the LDM autoencoder and the use of limited bitrate, we compare PerCo to the LDM encoder/decoder reconstruction and to PerCo trained without quantization in Tab. 1 (right). We observe that the LDM autoencoder significantly improves all metrics, and that PerCo trained without quantization obtains roughly the same performance as the autoencoder. This shows that PerCo is a strong conditional model, and that its performance is mainly bounded by the quantization module which throttles the bitrate of the image encoding, but could also lead to suboptimal training through noisy gradients.
@@
Number of timesteps as a tradeoff between perception and distortion. In Figure 8, we show how FID, LPIPS, MS-SSIM and CLIP scores vary when we change the number of denoising steps. We do the evaluation for seven different bitrates, ranging from 0.0032 to 0.1263 bpp. First, we notice that for higher bitrates, the metrics are quite constant across the number of timesteps, which means that the model has less uncertainty about the image to reconstruct and obtains optimal performance after only few denoising steps. Secondly, we observe a tradeoff between distortion and realism for lower bitrates. Indeed, when the number of timesteps is small, FID is worse, but MS-SSIM is optimal, while this phenomenon reverses when increasing the number of denoising steps.
@@ 5 Conclusion
Limitations. In this study we focused on medium sized images up to $768\!\times\!512$ resolution, similar to prior work on diffusion-based methods (Theis et al., 2022; Lei et al., 2023) that evaluated on $64\!\times\!64$ and $512\!\times\!512$ images, respectively. Extension to higher resolutions can possibly be achieved using a patch-based approach, see e.g. Hoogeboom et al. (2023).
PerCo exhibits somewhat poorer reconstruction performance in terms MS-SSIM, PSNR and LPIPS (the latter for rates $>$0.01), than existing approaches. This seems at least in part to be due to limitations of the LDM autoencoder, see Sec. B.2. We leave detailed study of these two points to future work.
@@ Appendix A Experimental details
Evaluation of third-party models.
Where possible we reported evaluation metrics reported in the original papers. For Text-Sketch Lei et al. (2023), we used the publicly released checkpoints, and used them for evaluation with the provided code. Note that Text-Sketch resized all images at resolution 512 for evaluation on the DIV2K and CLIC 2021 datasets, which makes the comparison with other baselines not immediate. We prefer to evaluate on original images instead, and for that purpose used the Kodak dataset which is commonly used to evaluate image compression models, and contains images of $512\!\times\!768$ pixels. We also evaluate on the MS-COCO 2014 validation set, which is less commonly used, but which offers ground truth semantic segmentation maps and descriptive captions which enable evaluation with the semantics oriented mIoU and CLIP score metrics.
@@
To finetune our latent diffusion model we used a grid of 50 timesteps. At inference time, we use 20 denoising steps to obtain good performance for the lowest bitrates. For bitrates of 0.05 and higher, we use five denoising steps as this is enough to obtain optimal performance, see Fig. 8 from the main paper. Besides, we sample images with DDIM deterministic sampling schedule, i.e. by fixing $\sigma_{t}=0$ for all $t\in[0,T]$, where $T$ is the number of denoising steps, see Song et al. (2021).
When testing different textual conditioning modalities (either BLIP or IDEFICS) in our ablation study in Fig. 7, the conditioning used at inference time is also used when training the models. When using human-specified captions, we use PerCo trained on BLIP captions, as there are no ground-truth captions for the OpenImagesv6 dataset we used to train our model.
@@
Details on the bitrates used to evaluate our models.
To achieve a specific target bitrate, we take into account the average bitrates required to encode the textual conditioning, and then set the spatial grid size and vocabulary size of the local conditioning accordingly. More specifically, for COCO at resolution $512\!\times\!512$, the average textual bpp for BLIP is 0.00219, while it is 0.00622 for IDEFICS, and 0.00271 for ground truth captions. For Kodak at resolution $768\!\times\!512$, it is 0.00141 with BLIP and 0.00415 bpp with IDEFICS. In the experiments, we select the spatial size and codebook dimension of the quantization module to adjust the spatial encoding bitrate depending on the targeted total bitrate. We found that codebooks with less than 64 elements were detrimental to the training, so we only used larger codebooks. Table 2 lists the bpp for the spatial coding for the configurations we used to proeuce the results in our experiments. For the ablation study in Fig. 7 we used other combinations of spatial grid size and vocabulary size for the tested variants to approximately match the bitrates of our default PerCo model.
Decoding time.

### v2

@@ 4.1 Experimental setup
Implementation details. We base our model off a text-conditioned latent diffusion model, with latent space provided by a convolutional autoencoder and diffusion model with an architecture similar to GLIDE (Nichol et al., 2022). The model is pre-trained on a proprietary dataset consisting of around 300M image-caption pairs. We also experiment with an image-conditioned model. We use BLIP-2 (Li et al., 2023) to obtain image captions for conditioning the decoder, and limit the maximum caption length to 32 tokens. The hyper encoder consists of nine residual convolutional blocks, and includes a number of downsampling layers depending on the spatial map resolution selected for the target bitrate. In total, the hyper encoder contains between 4M and 8M parameters. During all our experiments, the autoencoder weights are frozen. We train the hyper encoder and finetune the diffusion model on OpenImages (Kuznetsova et al., 2020), similar to Lei et al. (2023) and Muckley et al. (2023). We use random $512\!\times\!512$ crops, and instead of finetuning the full U-Net we found it beneficial to only finetune the linear layers present of the diffusion model, representing around 15% of all weights.
Datasets. For evaluation, we use the Kodak dataset (Franzen, 1999) as well as MS-COCO 30k. On COCO we evaluate at resolution $256\!\times\!256$ by selecting the same images from the 2014 validation set (Lin et al., 2014) as Hoogeboom et al. (2023) and Agustsson et al. (2023). We evaluate at resolution $512\!\times\!512$ on the 2017 training set (Caesar et al., 2018), which is the same resolution used for evaluation by Lei et al. (2023), and use captions and label maps for some metrics.
Metrics. To quantify image quality we use FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018), which match feature distributions between sets of original images and their reconstructions. We measure distortion using MS-SSIM (Wang et al., 2003) and LPIPS (Zhang et al., 2018). In Sec. B.1 we also report PSNR, but find these low-level distortion metrics to be less meaningful for low rates, see Figure 15. Therefore we include other more semantic-level metrics. We compute the CLIP score to measure global alignment of reconstructed samples with ground truth captions (Hessel et al., 2021). Unless otherwise specified, we condition our model on captions generated by BLIP2, which are different from the ground-truth captions used to compute the CLIP score. Second, we adopt the mean Intersection over Union (mIoU) metric to measure the local semantic preservation, see e.g. Sushko et al. (2022). In this case, we pass reconstructed images through a pre-trained semantic segmentation model, and compare the predicted semantic maps to the ground truth ones. We take the pretrained ViT-Adapter segmentation network (Chen et al., 2023), which was trained on the COCO train split. For Kodak we only report LPIPS and MS-SSIM, as the dataset is too small (24 images) to reliably compute FID and KID, and does not come with semantic annotations.
Baselines. We include representative codecs from several families. VTM is a state-of-the-art hand-crafted codec used in the VVC video codec (vtm, ). Among neural compressors using adversarial losses, we include MS-ILLM (Muckley et al., 2023) which improves over HiFiC (Mentzer et al., 2020) using an adversarial loss based on multi-class classification. For HiFiC we include their model with the lowest available bitrate, and add MS-PatchGAN, which is a close reproduction of HiFiC by Muckley et al. (2023). We compare to multi-realism approach of Agustsson et al. (2023), at their lowest bitrate with the best realism. Text-Sketch (Lei et al., 2023) is the only prior work we are aware of that evaluates image compression at bitrates below 0.01, it is based on text-conditioned diffusion models. We include two other recent diffusion-based approaches: DIRAC (Ghouse et al., 2023) and HFD/DDPM (Hoogeboom et al., 2023).
@@ 4.2 Main results
Figure 4: Comparisons on COCO at $256\!\times\!256$ resolution.
Comparison to state-of-the-art methods. In Figure 3 we compare our results to state-of-the-art codecs. We observe that PerCo (ours) yields significantly lower (better) FID and KID compared to other approaches at lower bitrates ($<$0.04 bpp), and that our FID and KID curves are much flatter, indicating a decoupling of realism and bitrate that is not seen for other methods. For the semantics-related CLIP and mIoU metrics we also obtain consistent improvements over all other methods, in particular at low rates. For LPIPS, PerCo is better than Text-Sketch and VTM, but somewhat worse than other approaches for rates $>$0.01 bpp. Similarly, for MS-SSIM we improve over Text-Sketch, but are worse than other methods. However, as we illustrate in figures 11 and 15 in Sec. B.2, such similarity metrics are not necessarily meaningful at low bitrates, as they do not capture realism. In Figure 4 we compare to HFD/DDPM (using their lowest bitrates: 0.107 and 0.175 bpp) on FID on MS-COCO 30k at $256\!\times\!256$ resolution, as used in their paper. We also include MS-ILLM and the multi-realism approach of Agustsson et al. (2023). We find that PerCo yields best FID for both rates.
@@
Qualitative comparisons. We provide qualitative comparisons of PerCo to baselines in Figure 1 and Figure 5. PerCo yields reconstructions of higher quality with less artefacts, more faithful colors, and realistic details. VTM and MS-ILLM reconstructions are blurry and/or blocky, while Text-Sketch strongly deviates from the originals in color. When using text-only conditioning for Text-Sketch (PIC, 0.0025 bpp), it is not able to achieve spatial alignment, by lack of local image encoding.
4.3 Ablations
Diversity in reconstructions. With our model, reconstructions are sampled from a conditional diffusion model. Ideally, the conditional distribution concentrates all mass perfectly on the original input image. In practice, this does not happen due to limited model capacity, training data, and bitrate. In particular, we expect that at lower bitrates —where there is less information about the original sample— the conditional distribution would have more variance, and we expect to observe more diversity among the samples. The reconstruction at two bitrates in Figure 6 confirm this.
Conditioning modalities. In Figure 7, we evaluate PerCo using each conditioning modality separately (textual or spatial), to analyze the contribution of each. We compare BLIP-2 (Li et al., 2023) to generate captions (default) to using IDEFICS (Laurançon et al., 2023) which produces more detailed descriptions, and ground-truth (GT) captions. When not using descriptive captions, we consider conditioning on a learned textual embedding that is constant across images. We also tried using an empty text or on a fixed generic text —“A high quality photograph.”— for all images, and observed similar or slightly worse performance as the learned text embedding setting.
The model using the constant text embedding (PerCo w/ only Spatial Encoding) performs significantly worse than the other models in terms of FID and CLIP score. For the LPIPS and mIoU metrics, which rely on the alignment of the original and reconstructed images, the results are comparable to other models. This can be attributed to the fact that accurate alignment is primarily achieved through local feature conditioning, which remains unchanged in this model. Consistently, we also notice that when using textual condition only (blue, orange, and red square), the performance in terms of LPIPS and mIoU is worse than for the models that use spatial conditioning, while the CLIP scores are in line with the models using both textual and spatial conditioning. Both types of conditioning contribute to improving FID, and the text-only conditioned models perform in between the models with only spatial conditioning and those with both types of conditioning.
Between models using captions from BLIP (blue curves) and IDEFICS (orange curves) in combination with spatial conditioning, we observe similar FID scores and somewhat better CLIP scores for the more detailed IDEFICS captions. In terms of LPIPS and mIoU we find that the models with BLIP captions perform better, which is largely due to bitrate required to encode the longer captions: leaving out the textual encoding cost, both models perform very similarly. Overall, using ground-truth captions (GT, red curves) yields similar results as using BLIP, while BLIP and IDEFICS yield slightly better mIoU and CLIP scores. As COCO contains five GT captions per image, we ensure that when conditioning on a GT caption, a different one is used to compute the CLIP score.
@@ 4.3 Ablations
Classifier-free guidance. In Table 1 (left), we compare three options regarding classifier-free guidance (CFG) in our model (i) not using CFG, (ii) applying CFG on the text and spatial feature maps, i.e. contrasting full conditioning to no conditioning (we use an constant learned text embedding, and fixed all-zeros feature maps for local conditioning), and (iii) applying CFG to text-conditioning only, i.e. always conditioning on the spatial feature map. We find that best results are obtained when CFG is applied to textual conditioning only, which yields significantly better FID and mIoU metrics.
Impact of autoencoder and quantization. To assess to which extent performance is bounded by the LDM autoencoder or the use of limited bitrate, we compare PerCo to the LDM encoder/decoder reconstruction and to PerCo trained without quantization in Table 1 (right). We observe that the LDM autoencoder significantly improves all metrics, and that PerCo trained without quantization obtains roughly the same performance as the autoencoder. This shows that PerCo is a strong conditional model, and that its performance is mainly bounded by the quantization module which throttles the bitrate of the image encoding, but could also lead to suboptimal training through noisy gradients.
@@
Number of timesteps as a tradeoff between perception and distortion. In Figure 8, we show how FID, LPIPS, MS-SSIM and CLIP scores evolve along the denoising steps. We do the evaluation for seven different bitrates, ranging from 0.0032 to 0.1263 bpp. First, we notice that for higher bitrates, the metrics are quite constant across the number of timesteps, which means that the model has less uncertainty about the image to reconstruct and obtains optimal performance after only few denoising steps. Secondly, we observe a tradeoff between distortion and realism for lower bitrates. Indeed, when the number of timesteps is small, FID is worse, but MS-SSIM is optimal, while this phenomenon reverses when increasing the number of denoising steps.
Conditioning on global image features. Besides the text-conditioned model described above and used in most experiments (PerCo), we also experimented with a model conditioned on image embeddings extracted from an image backbone network. While we use lossless Lempel-Ziv coding to compress the captions for the text-conditioned model, we use product quantization (Jégou et al., 2011) for lossy compression of the image embeddings. We divide the embeddings into $M$ subvectors and quantize each subvector using a separate codebook with $V$ elements. We empirically found that we achieve similar performance compared to training with non-quantized image embeddings with $M=16$ and $V=1024$. For $512\!\times\!512$ images, the bitrate obtained for the image embeddings is 0.00061 bpp, which is 3.5$\times$ lower than the bitrate of 0.00219 bpp for BLIP image captions. From the results in Figure 3, we observe that the model with global image embeddings obtains competitive results, but has slightly worse pairwise image reconstruction metrics at high bitrates in terms of LPIPS and MS-SSIM on Kodak and MS-COCO 30k.
@@ 5 Conclusion
Limitations. In this study we focused on medium sized images up to $768\!\times\!512$ resolution, similar to prior work on diffusion-based methods (Theis et al., 2022; Lei et al., 2023) that evaluated on $64\!\times\!64$ and $512\!\times\!512$ images, respectively. Extension to higher resolutions can possibly be achieved using a patch-based approach, see e.g. Hoogeboom et al. (2023). PerCo exhibits somewhat poorer reconstruction performance in terms MS-SSIM, PSNR and LPIPS (the latter for rates $>$0.01), than existing approaches. This seems at least in part to be due to limitations of the LDM autoencoder, see Sec. B.2. It is also probably explained by the tradeoff existing between distortion and perception (Blau & Michaeli, 2019). We leave detailed study of these two points to future work.
@@ Appendix A Experimental details
Evaluation of third-party models. Where possible we reported evaluation metrics reported in the original papers. For Text-Sketch Lei et al. (2023), we used the publicly released checkpoints, and used them for evaluation with the provided code. Note that Text-Sketch resized all images at resolution 512 for evaluation on the DIV2K and CLIC 2021 datasets, which makes the comparison with other baselines not immediate. We prefer to evaluate on original images instead, and for that purpose used the Kodak dataset which is commonly used to evaluate image compression models, and contains images of $512\!\times\!768$ pixels. We also evaluate on the MS-COCO 2014 validation set, which is less commonly used, but which offers ground truth semantic segmentation maps and descriptive captions which enable evaluation with the semantics oriented mIoU and CLIP score metrics.
@@
To finetune our latent diffusion model we used a grid of 50 timesteps. At inference time, we use 20 denoising steps to obtain good performance for the lowest bitrates. For bitrates of 0.05 and higher, we use five denoising steps as this is enough to obtain optimal performance, see Figure 8 from the main paper. Besides, we sample images with DDIM deterministic sampling schedule, i.e. by fixing $\sigma_{t}=0$ for all $t\in[0,T]$, where $T$ is the number of denoising steps, see Song et al. (2021).
When testing different textual conditioning modalities (either BLIP or IDEFICS) in our ablation study in Figure 7, the conditioning used at inference time is also used when training the models. When using human-specified captions, we use PerCo trained on BLIP captions, as there are no ground-truth captions for the OpenImagesv6 dataset we used to train our model.
@@
Details on the bitrates used to evaluate our models. To achieve a specific target bitrate, we take into account the average bitrates required to encode the textual conditioning, and then set the spatial grid size and vocabulary size of the local conditioning accordingly. More specifically, for COCO at resolution $512\!\times\!512$, the average textual bpp for BLIP is 0.00219, while it is 0.00622 for IDEFICS, and 0.00271 for ground truth captions. For Kodak at resolution $768\!\times\!512$, it is 0.00141 with BLIP and 0.00415 bpp with IDEFICS. In the experiments, we select the spatial size and codebook dimension of the quantization module to adjust the spatial encoding bitrate depending on the targeted total bitrate. We found that codebooks with less than 64 elements were detrimental to the training, so we only used larger codebooks. Table 2 lists the bpp for the spatial coding for the configurations we used to proeuce the results in our experiments. For the ablation study in Figure 7 we used other combinations of spatial grid size and vocabulary size for the tested variants to approximately match the bitrates of our default PerCo model.
Inference Speed. We run benchmarking on A100 GPUs using 20 denoising steps, and 5 denoising steps for bitrates higher than 0.05 bits per pixel. We compute the runtime for all the Kodak images at resolution $512\times 768$ and provide the average time in seconds with its standard deviation in Table 3. The encoding complexity of PerCo is quite similar to adversarial-based compression methods such as MS-ILLM and HiFiC (which has the same architecture as MS-ILLM). Encoding with Text-Sketch (PIC/PICS) is orders of magnitude slower because it is optimizing a text embedding for each image. The decoding complexity of PerCo using 20 denoising steps (bitrates below 0.05 bpp) is comparable to that of Text-Sketch (PIC) and about 8 times faster than Text-Sketch (PICS), because PICS is conditioning on sketches which takes time to decompress. When using only five denoising steps, the decoding complexity drops rouhgly by a factor four, showing that the decoding is dominated by evaluations of the denoising network. Compared to HiFiC and MS-ILLM, the PerCo decoder using 5 denoising steps is about eight times slower, and when using 20 steps this becomes a factor 32. Please note that the running time of VTM is computed on CPU, so it may explain the longer runtime. We could not evaluate the diffusion-based DIRAC and HFD/DDPM models as there is no public codebase available, but we found some evaluations in DIRAC paper about decoding time. They report a decoding time which is about four times slower than what they report for HiFiC, which also shows that diffusion-based methods are slower than adversarial-based ones.
Model Encoding Speed (in sec.) Decoding Speed (in sec.)
VTM 16.892 $\pm$ 7.574 0.135 $\pm$ 0.002
MS-ILLM 0.084 $\pm$ 0.021 0.080 $\pm$ 0.007
Text-Sketch (PIC) 163.070 $\pm$ 0.380 2.725 $\pm$ 0.012
Text-Sketch (PICS) 190.231 $\pm$ 2.476 19.288 $\pm$ 0.251
PerCo - 5 denoising steps 0.080 $\pm$ 0.018 0.665 $\pm$ 0.009
PerCo - 20 denoising steps 0.080 $\pm$ 0.018 2.551 $\pm$ 0.018
Table 3: Encoding and decoding Speed (in seconds).